### v2

@@ Abstract
Interacting particle systems in a finite volume in equilibrium are often described by a grand canonical ensemble induced by the corresponding Hamiltonian, i.e. a finite volume Gibbs measure. However, in practice, directly measuring this Hamiltonian is not possible, as such, methods need to be developed to calculate the Hamiltonian potentials from measurable data. In this work, we give an expansion of the chemical potential in terms of the correlation functions of such a system in the thermodynamic limit. This is a justification of a formal approach of Nettleton and Green from the 50’s, that can be seen as an inverse cluster expansion.
@@ 1 Introduction
In classical equilibrium statistical physics, interacting particle systems are often described via so-called Gibbs measures, either in a finite or infinite volume. A possible arrangement of particles (also called configuration) has a higher probability when it is energetically favourable. This energy gets calculated via a Hamiltonian that depends on the chemical potential and the interaction potentials associated to the particles. However, when working with actual data, these potentials are impossible to measure. What instead can be measured or extrapolated from simulation data are the so-called correlation functions of the underlying Gibbs measure. This gives rise to the inverse problem: Given the correlation functions, determine the chemical potential and the interaction potentials.
@@
For the pair interaction case the inverse problem was investigated in [14] in the case of the finite volume and in [17] and [12] these results were extended upon and made rigorous in the thermodynamic limit, i.e. in infinite volume. However, all these results do not give a constructive way of recovering the interaction. Estimation methods for the pair interaction were developed by Takacs and Fiksel in [29] and [9] and have since been expanded upon for more general Hamiltonians in e.g. [6], another approach based on variational methods was developed in [2], however, it cannot be used to estimate the chemical potential. A first result of an inverse expansion of the chemical potential is found in [19] where the expansion is given in terms of the one-point correlation function (the density) and the Mayer function, see also [15] for an extension of this result to the non-homogeneous case. For this formula however prior knowledge of the pair interaction is required. The goal of this paper is to find an expansion of the chemical potential, that solely depends on functions derived from the correlation functions of the Gibbs measure.
The idea we use goes back to [21], where an expansion of this type is used in finite volume. However, the ansatz used in these works constructs multi-body interactions and thus showing convergence of these inversion formulas cannot be done using the classical tools of cluster expansions.
In this work, we will use a different approach to prove convergence for the explicit inversion formula for the chemical potential by using an exponential representation formula. We then use Bell-polynomials to find bounds on the appearing coefficients, which have been used for a similar scope independently in [22] to get bounds on the virial coefficients. In fact, we can show that this formula converges in both finite and infinite volume. In particular, as we only make assumptions on the correlation functions and not the Hamiltonian, the expansion does not only hold for two-body interactions and even holds in some cases for multi-body interactions of arbitrary order.
@@ 2 Setting
equipped with the $\sigma$-algebra $\mathcal{F}:=\sigma\left(\{N(\cdot\cap\Delta)=m\}\mid m\in{\mathord{\mathbb{N}}}_{0},\,\,\Delta\subset{\mathord{\mathbb{R}}}^{d}\text{ bounded}\right)$. Here $N(\eta)=\#\eta$ is the number of elements of $\eta\subset{\mathord{\mathbb{R}}}^{d}$. We denote by $\Gamma_{0}=\{\gamma\in\Gamma\mid N(\gamma)<\infty\}$ the set of finite configurations. Such a point process is a Gibbs measure for some chemical potential $\mu\in{\mathord{\mathbb{R}}}$ and a Hamiltonian $H\colon\Gamma\to{\mathord{\mathbb{R}}}\cup\{\infty\}$, if it is supported on a set of tempered configurations, see e.g. [8] for a discussion, and satisfies the Ruelle-equation, see [26] (herein called system of equilibrium equations), namely if, for every non-negative function $G\colon\Gamma\to[0,\infty]$ and every bounded set $\Lambda\subset{\mathord{\mathbb{R}}}^{d}$, there holds
@@
where $\Lambda^{c}={\mathord{\mathbb{R}}}^{d}\backslash\Lambda$, $\Gamma_{\Lambda^{c}}=\{\eta\in\Gamma\mid\eta\subset\Lambda^{c}\}$ and
$W(\{\textit{{x}}_{n}\}\mid\eta)=H(\{\textit{{x}}_{n}\}\cup\eta)-H(\{\textit{{x}}_{n}\})-H(\eta),\qquad\textit{{x}}_{n}\in\Lambda^{n},\,\eta\in\Gamma_{\Lambda^{c}}$
is the interaction between the part of the particles inside $\Lambda$ and those on the outside, with the convention that we set $W(\{\textit{{x}}_{n}\}\mid\eta)=+\infty$ if $H(\{\textit{{x}}_{n}\}\cup\eta)=+\infty$. We write ${\mathord{\mathbb{P}}}$ is a $(\mu,H)$-Gibbs measure. We assume that the Hamiltonian $H$ is translation invariant, i.e. for every $x\in{\mathord{\mathbb{R}}}^{d}$ we have that $H(\eta)=H(\tau_{x}(\eta))$ where $\tau_{x}(\eta):=\{y-x\mid y\in\eta\}$, that $H$ is hereditary, i.e. that $H(\eta\cup\{x\})=+\infty$ for all $x\in{\mathord{\mathbb{R}}}^{d}$ whenever $H(\eta)=+\infty$ and that $H$ does not contain any contribution of a self-interaction of particles, i.e. $H(\{x\})=0$ for all $x\in{\mathord{\mathbb{R}}}^{d}$. The last assumption we make on $H$ is, that $H$ is stable, i.e. there is a constant $B>0$ such that
@@
In this case, under some additional technical assumptions, e.g. finite range of the interaction or the Hamiltonian consists of a regular pair interaction, there is at least one translation invariant $(\mu,H)$-Gibbs measure ${\mathord{\mathbb{P}}}$ associated to these parameters, cf. [8]. In the sequel, we will assume that ${\mathord{\mathbb{P}}}$ is such a translation invariant Gibbs measure. A family of functions $(j_{\Lambda}^{(n)})_{n\in\geq 0,\,\Lambda\subset{\mathord{\mathbb{R}}}^{d}\text{bounded}}$ are called the finite volume densities of ${\mathord{\mathbb{P}}}$, or Janossy densities, if for every bounded set $\Lambda\subset{\mathord{\mathbb{R}}}^{d}$ and every nonnegative function $G\colon\Gamma\to[0,\infty]$ there holds
@@
From the Ruelle-equation it also follows that exist for every $(\mu,H)$-Gibbs measure $/P$ and for fixed $\Lambda\subset{\mathord{\mathbb{R}}}^{d}$ they are given by
@@
for any $\Lambda$ such that $\textit{{x}}_{n}\in\Lambda^{n}$. Note that we write $j_{\Lambda}^{(n+k)}(\textit{{x}}_{n},\textit{{y}}_{k})=j_{\Lambda}^{(n+k)}(x_{1},\dots,x_{n},y_{1},\dots,y_{k})$ for simplicity and that the Janossy densities contain the averaged contributions of the whole space by (1) and thus the right-hand side above is independent of $\Lambda$, if $\textit{{x}}_{n}\in\Lambda^{n}.$ In particular, if ${\mathord{\mathbb{P}}}$ is translation invariant, then $\rho^{(1)}=\rho$ is constant. Under the additional assumption that ${\mathord{\mathbb{P}}}$ satisfies a so-called Ruelle condition, i.e. there is a $\xi>0$ such that
@@
see e.g. [26]. The idea of this work is to find an explicit expansion of $\log j_{\Lambda}^{(1)}$, using the identity (2), in terms that only depend on the correlation functions $(\rho^{(n)})_{n\geq 0}$. From this, we can then obtain an explicit formula for the chemical potential $\mu$. The reason is as follows:
@@
Now it is well-known that both sides of the above equation diverge to $\infty$- ∞ for $\Lambda\nearrow{\mathord{\mathbb{R}}}^{d}$ (here $\Lambda\nearrow{\mathord{\mathbb{R}}}^{d}$ means that we take an increasing sequence of bounded sets $\Lambda$ such that eventually every bounded set of ${\mathord{\mathbb{R}}}^{d}$ is contained in them), but in the sequel, we will show that for the examples we look at there holds
@@ 3 Main result
The Janossy densities of ${\mathord{\mathbb{P}}}$ satisfy: for every compact set $\Delta\subset{\mathord{\mathbb{R}}}^{d}$ there holds
$\log j_{\Lambda}^{(1)}(x)-\log j_{\Lambda}^{(0)}\to\mu,$ (7)
uniformly on $\Delta$, as $\Lambda\nearrow{\mathord{\mathbb{R}}}^{d}$.
@@
$\sup_{x\in{\mathord{\mathbb{R}}}^{d}}\int_{\Lambda^{n}}\left|{\rho}_{T}^{(1+n)}(x,\textit{{y}}_{n})\right|\,\mathrm{d}\textit{{y}}_{n}\leq n!Dq^{n},$ (8)
@@
Let ${\mathord{\mathbb{P}}}$ be a $(\mu,H)$-Gibbs measure that satisfies assumptions A and B. If $q<q_{0}$, where
@@ 4 Examples
We will now take a look at three cases in which the assumptions of Theorem 3.1 are satisfied. First, we will look at a classic superstable pair interaction, then at a Hamiltonian consisting of multi-body potentials of arbitrary order, which are non-negative and of finite range, in both settings there always exists at least one translation invariant $(\mu,H)$-Gibbs measure, cf. [8]. Finally, we will look at the Kirkwood-closure process from [18] which can also be shown to be a Gibbs point process for a Hamiltonian consisting of multi-body potentials of every order.
4.1 Superstable pair interaction
We take a look at the same setting as in Remark 5.1. Let ${\mathord{\mathbb{P}}}$ be such a $(\mu,u)$-Gibbs measure, where $u$ is a superstable pair interaction. From Theorem 5.5 in [26], it follows immediately that ${\mathord{\mathbb{P}}}$ satisfies assumption A. Furthermore, we note that, by Remark 5.1, ${\mathord{\mathbb{P}}}$ satisfies assumption B. We thus get the easy result:
@@ 4.1 Superstable pair interaction
Let ${\mathord{\mathbb{P}}}$ be a $(\mu,u)$-Gibbs measure with a superstable pair interaction $u$, if
@@
Note that the constants appearing in the upper bound on $\mu$ are discussed in Remark 5.1.
4.2 Non-negative multi-body interaction of finite range
We say a Hamiltonian has finite range, if there exists an $R>0$ such that, whenever $|x_{0}-y|>R$ there holds
@@ 4.2 Non-negative multi-body interaction of finite range
for every set of points $\{\textit{{x}}_{n}\}\subset{\mathord{\mathbb{R}}}^{d}$. In particular, this means that $W(\{x_{0}\}\mid\eta)=0$, for any $\eta\in\Gamma$ with $\textnormal{dist}(\{x_{0}\},\eta)>R$. As mentioned, in this case there always exists at least one $(\mu,H)$-Gibbs measure, see e.g. [8]. Furthermore, ${\mathord{\mathbb{P}}}$ satisfies a Ruelle condition with $\xi=e^{\mu}$.
@@
$j_{\Lambda}^{(1)}(x)=e^{\mu}\int_{\Gamma_{\Lambda^{c}}}\,\,\mathrm{d}{\mathord{\mathbb{P}}}(\eta)=e^{\mu}j_{\Lambda}^{(0)}.$
@@
It was shown by Moraal in [20] that if $H$ is additionally non-negative, then the multi-body version of the Kirkwood-Salsburg operator is bounded and there exists a solution to the multi-body version of the Kirkwood-Salsburg equation. Using the same techniques as in the pair interaction case, see Section 4.4.6 of [25], it then follows that ${\mathord{\mathbb{P}}}$ satisfies assumption B.
@@
In [27] Skrypnik extended the results of Moraal to the case that $H$ consists of the sum of a very particular non-negative multi-body interaction of finite range and a superstable pair interaction. In this case, it can also be shown that assumption A is fulfilled.

### v3

@@ Abstract
Interacting particle systems in a finite-volume in equilibrium are often described by a grand-canonical ensemble induced by the corresponding Hamiltonian, i.e. a finite-volume Gibbs measure. However, in practice, directly measuring this Hamiltonian is not possible, as such, methods need to be developed to calculate the Hamiltonian potentials from measurable data. In this work, we give an expansion of the chemical potential in terms of the correlation functions of such a system in the thermodynamic limit. This is a justification of a formal approach of Nettleton and Green from the 50’s, that can be seen as an inverse cluster expansion.
@@ 1 Introduction
In classical equilibrium statistical physics, interacting particle systems are often described via so-called Gibbs measures, either in a finite or infinite-volume. A possible arrangement of particles (also called configuration) has a higher probability when it is energetically favorable. This energy gets calculated via a Hamiltonian that depends on the chemical potential and the interaction potentials associated to the particles. However, when working with actual data, these potentials are impossible to measure. What instead can be measured or extrapolated from simulation data are the so-called correlation functions of the underlying Gibbs measure. This gives rise to the inverse problem: Given the correlation functions, determine the chemical potential and the interaction potentials.
@@
For the pair-interaction case the inverse problem was investigated in [14] in the case of the finite-volume and in [17] and [12] these results were extended upon and made rigorous in the thermodynamic limit, i.e. in infinite-volume. However, all these results do not give a constructive way of recovering the interaction. Estimation methods for the pair-interaction were developed by Takacs and Fiksel in [29] and [9] and have since been expanded upon for more general Hamiltonians in e.g. [6], another approach based on variational methods was developed in [2], however, it cannot be used to estimate the chemical potential. A first result of an inverse expansion of the chemical potential is found in [19] where the expansion is given in terms of the one-point correlation function (the density) and the Mayer function, see also [15] for an extension of this result to the non-homogeneous case. For this formula however prior knowledge of the pair-interaction is required. The goal of this paper is to find an expansion of the chemical potential, that solely depends on functions derived from the correlation functions of the Gibbs measure.
The idea we use goes back to [21], where an expansion of this type is used in finite-volume. However, the ansatz used in these works constructs multi-body interactions and thus showing convergence of these inversion formulas cannot be done using the classical tools of cluster expansions.
In this work, we will use a different approach to prove convergence for the explicit inversion formula for the chemical potential by using an exponential representation formula. We then use Bell-polynomials to find bounds on the appearing coefficients, which have been used for a similar scope independently in [22] to get bounds on the virial coefficients. In fact, we can show that this formula converges in both finite and infinite-volume. In particular, as we only make very mild assumptions on the Hamiltonian, which include many classic models, and some integrability conditions on the correlation functions this expansion holds not only for pair-interactions, but also in some cases of multi-body interactions of arbitrary order.
The outline is as follows: In Section 2 we introduce the general setting of Gibbs measures we work with and introduce the main idea of the expansion. Then, we formulate in Section 3 our main assumptions as well as the main result and give a few examples in which these assumptions hold in Section 4. The last two Sections are devoted to the proof of the main result, in Section 5 we start by proving a result about the zero-point density, which then gets generalized in Section 6 to the one-point density to obtain the expansion for the chemical potential.
@@ 2 Setting
equipped with the $\sigma$-algebra $\mathcal{F}:=\sigma\left(\{N(\cdot\cap\Delta)=m\}\mid m\in{\mathord{\mathbb{N}}}_{0},\,\,\Delta\subset{\mathord{\mathbb{R}}}^{d}\text{ bounded}\right)$. Here $N(\eta)=\#\eta$ is the number of elements of $\eta\subset{\mathord{\mathbb{R}}}^{d}$. We denote by $\Gamma_{0}=\{\gamma\in\Gamma\mid N(\gamma)<\infty\}$ the set of finite configurations. Such a point process is a Gibbs measure for some chemical potential $\mu\in{\mathord{\mathbb{R}}}$ and a Hamiltonian $H\colon\Gamma\to{\mathord{\mathbb{R}}}\cup\{\infty\}$, we write ${\mathord{\mathbb{P}}}$ is a $(\mu,H)$-Gibbs measure, if it is supported on a set of tempered configurations, see e.g. [8] for a discussion, and satisfies the Ruelle-equation, see [26] (herein called system of equilibrium equations), namely if, for every non-negative function $G\colon\Gamma\to[0,\infty]$ and every bounded set $\Lambda\subset{\mathord{\mathbb{R}}}^{d}$, there holds
@@
where $\Lambda^{c}={\mathord{\mathbb{R}}}^{d}\backslash\Lambda$, $\Gamma_{\Lambda^{c}}=\{\eta\in\Gamma\mid\eta\subset\Lambda^{c}\}$ and $W(\{\textit{{x}}_{n}\}\mid\eta)$ is the interaction between the part of the particles inside $\Lambda$ and those on the outside. If $H$ has finite-range $W$ can be defined as
$W(\{\textit{{x}}_{n}\}\mid\eta):=H(\{\textit{{x}}_{n}\}\cup\eta)-H(\{\textit{{x}}_{n}\})-H(\eta),\qquad\textit{{x}}_{n}\in\Lambda^{n},\,\eta\in\Gamma_{\Lambda^{c}}$
with the convention that we set $W(\{\textit{{x}}_{n}\}\mid\eta)=+\infty$ if $H(\{\textit{{x}}_{n}\}\cup\eta)=+\infty$. For general Hamiltonians it is possible to extend this definition, cf. [8]. However, the specific definition of $W$ is not important for this work. We also assume that the Hamiltonian $H$ is translation-invariant, i.e. for every $x\in{\mathord{\mathbb{R}}}^{d}$ we have that $H(\eta)=H(\tau_{x}(\eta))$ where $\tau_{x}(\eta):=\{y-x\mid y\in\eta\}$, that $H$ is hereditary, i.e. that $H(\eta\cup\{x\})=+\infty$ for all $x\in{\mathord{\mathbb{R}}}^{d}$ whenever $H(\eta)=+\infty$ and that $H$ does not contain any contribution of a self-interaction of particles, i.e. $H(\{x\})=0$ for all $x\in{\mathord{\mathbb{R}}}^{d}$. The last assumption we make on $H$ is, that $H$ is stable, i.e. there is a constant $B>0$ such that
@@
In this case, under some additional technical assumptions, e.g. finite-range of the interaction or the Hamiltonian consists of a regular pair-interaction, there is at least one translation-invariant $(\mu,H)$-Gibbs measure ${\mathord{\mathbb{P}}}$ associated to these parameters, cf. [8]. In the sequel, we will assume that ${\mathord{\mathbb{P}}}$ is such a translation-invariant Gibbs measure. A family of functions $(j_{\Lambda}^{(n)})_{n\geq 0,\,\Lambda\subset{\mathord{\mathbb{R}}}^{d}\text{ bounded}}$ are called the finite-volume densities of ${\mathord{\mathbb{P}}}$, or Janossy densities, if for every bounded set $\Lambda\subset{\mathord{\mathbb{R}}}^{d}$ and every non-negative function $G\colon\Gamma\to[0,\infty]$ there holds
@@
The Ruelle-equation characterizes the Janossy densities, they exist for every $(\mu,H)$-Gibbs measure ${\mathord{\mathbb{P}}}$ and for fixed $\Lambda\subset{\mathord{\mathbb{R}}}^{d}$ they are given by
@@
for any $\Lambda$ such that $\textit{{x}}_{n}\in\Lambda^{n}$. Note that we write $j_{\Lambda}^{(n+k)}(\textit{{x}}_{n},\textit{{y}}_{k})=j_{\Lambda}^{(n+k)}(x_{1},\dots,x_{n},y_{1},\dots,y_{k})$ for simplicity and that the Janossy densities contain the averaged contributions of the whole space by (1) and thus the right-hand side above is independent of $\Lambda$, if $\textit{{x}}_{n}\in\Lambda^{n}.$ In particular, if ${\mathord{\mathbb{P}}}$ is translation-invariant, then $\rho^{(1)}=\rho$ is constant. Under the additional assumption that ${\mathord{\mathbb{P}}}$ satisfies a so-called Ruelle-condition, i.e. there is a $\xi>0$ such that
@@
see e.g. [26].
Remark 2.1.
In experiments, when only having samples of a Gibbs measure, it is straightforward to estimate the correlation functions of the process without having to make prior assumptions on the Hamiltonian. Thus, the correlation functions are the experimentally easier to obtain starting object in contrast to the Janossy densities.
The idea of this work is to find an explicit expansion of $\log j_{\Lambda}^{(1)}$, using the identity (2), in terms that only depend on the correlation functions $(\rho^{(n)})_{n\geq 0}$. From this, we can then obtain an explicit formula for the chemical potential $\mu$. The reason is as follows:
@@
Now it is well-known that both sides of the above equation diverge to $\infty$- ∞ for $\Lambda\nearrow{\mathord{\mathbb{R}}}^{d}$. In the sequel we will make the following technical assumption, that can be interpreted as an implicit decay condition on the Hamiltonian and in particular is obvious for finite-range potentials,
@@ 3 Main result
Let $W$ be the interaction associated to the Hamiltonian $H$. We assume for any bounded set $\Delta\subset{\mathord{\mathbb{R}}}^{d}$ and every $(\mu,H)$-Gibbs measure ${\mathord{\mathbb{P}}}$ there holds
$\lim_{\Lambda\nearrow{\mathord{\mathbb{R}}}^{d}}\sup_{x\in\Delta}\left|\log\int_{\Gamma_{\Lambda^{c}}}\exp\Big{(}-W(\{x\}\mid\eta)\Big{)}\,\,\mathrm{d}{\mathord{\mathbb{P}}}(\eta)-\log\int_{\Gamma_{\Lambda^{c}}}\,\,\mathrm{d}{\mathord{\mathbb{P}}}(\eta)\right|=0.$
@@
$\sup_{x\in{\mathord{\mathbb{R}}}^{d}}\int_{\Lambda^{n}}\frac{1}{\rho}\left|{\rho}_{T}^{(1+n)}(x,\textit{{y}}_{n})\right|\,\mathrm{d}\textit{{y}}_{n}\leq n!Dq^{n},$ (7)
@@
Let ${\mathord{\mathbb{P}}}$ be a $(\mu,H)$-Gibbs measure that satisfies a Ruelle-conditions and Assumptions A and B. If $q<q_{0}$, where
@@
Remark 3.2.
The functions $(\widetilde{\rho}^{\,(1+k)}_{T})_{k\geq 1}$ take a similar role as the so-called Ursell functions in classical cluster expansions and thus the expansion above can be seen as a type of multi-body cluster expansion.
@@ 4 Examples
We will now take a look at three cases in which the assumptions of Theorem 3.1 are satisfied. First, we will look at a classic superstable pair-interaction, then at a Hamiltonian consisting of multi-body potentials of arbitrary order, which are non-negative and of finite-range, in both settings there always exists at least one translation-invariant $(\mu,H)$-Gibbs measure, cf. [8]. Finally, we will look at the Kirkwood-closure process from [18] which can also be shown to be a Gibbs point process for a Hamiltonian consisting of multi-body potentials of every order.
4.1 Superstable pair-interaction
Let us consider the case that $H$ consists only of a pair-interaction, i.e. a measurable even function $u\colon{\mathord{\mathbb{R}}}^{d}\to{\mathord{\mathbb{R}}}\cup\{+\infty\}$ such that $H(\{\textit{{x}}_{n}\})=\sum_{i<j}u(x_{i}-x_{j})$. If we further assume that $u$ is superstable, i.e. that there exists $r_{0}>0$ and decreasing positive functions $\varphi:(0,r_{0})\to{\mathord{\mathbb{R}}}^{+}_{0}$ and $\psi:[0,\infty)\to{\mathord{\mathbb{R}}}^{+}$ with
and
$u(x)$ $\,\geq\,\varphi(|x|)\qquad\textrm{ for }\ 0<|x|<r_{0}\,,$
$|u(x)|$ $\,\leq\,\psi(|x|)\qquad\textrm{ for }\ |x|\geq r_{0}\,.$
It is well-known, cf. Chapter 4 of [25] that ${\mathord{\mathbb{P}}}$ satisifies Assumption B if
$e^{\mu}=\frac{\bar{q}}{e^{2B+1}C(u)}$ (11)
for some $\bar{q}<1/2$, where $C(u):=\int_{{\mathord{\mathbb{R}}}^{d}}|e^{-u(x)}-1|\,\mathrm{d}x$ and $B=B(u)>0$ is the stability constant of $H$, with
$D=\frac{e^{\mu}e^{-2B}}{(1-\bar{q})}\frac{1-\bar{q}}{1-2\bar{q}}\qquad\text{ and }\qquad q=\frac{\bar{q}}{1-\bar{q}}.$
Let ${\mathord{\mathbb{P}}}$ be such a $(\mu,u)$-Gibbs measure. From Theorem 5.5 in [26], it follows immediately that ${\mathord{\mathbb{P}}}$ satisfies Assumption A. Therein, it is also shown that ${\mathord{\mathbb{P}}}$ satisfies a Ruelle-condition. We thus get the easy result:
@@ 4.1 Superstable pair-interaction
Let ${\mathord{\mathbb{P}}}$ be a $(\mu,u)$-Gibbs measure with a superstable pair-interaction $u$, if
@@
4.2 Non-negative multi-body interaction of finite-range
We say a Hamiltonian has finite-range, if there exists an $R>0$ such that, whenever $|x_{0}-y|>R$ there holds
@@ 4.2 Non-negative multi-body interaction of finite-range
for every set of points $\{\textit{{x}}_{n}\}\subset{\mathord{\mathbb{R}}}^{d}$. In particular, this means that $W(\{x_{0}\}\mid\eta)=0$, for any $\eta\in\Gamma$ with $\textnormal{dist}(\{x_{0}\},\eta)>R$. As mentioned, in this case there always exists at least one $(\mu,H)$-Gibbs measure, see e.g. [8]. Furthermore, ${\mathord{\mathbb{P}}}$ satisfies a Ruelle-condition with $\xi=e^{\mu}$.
@@
$j_{\Lambda}^{(1)}(x)=e^{\mu}\int_{\Gamma_{\Lambda^{c}}}\,\,\mathrm{d}{\mathord{\mathbb{P}}}(\eta).$
@@
It was shown by Moraal in [20] that if $H$ is additionally non-negative, then the multi-body version of the Kirkwood-Salsburg operator is bounded and there exists a solution to the multi-body version of the Kirkwood-Salsburg equation. Using the same techniques as in the pair-interaction case, see Section 4.4.6 of [25], it then follows that ${\mathord{\mathbb{P}}}$ satisfies Assumption B.
@@
In [27] Skrypnik extended the results of Moraal to the case that $H$ consists of the sum of a very particular non-negative multi-body interaction of finite-range and a superstable pair-interaction. In this case, it can also be shown that Assumption A is fulfilled.